Two bodies are projected from the same point in directions inclined at 60 and 30 to the horizontal . If they attain the same maximum height the ratio of their velocities of projection is

Grade:11

1 Answers

Abhiram
7 years ago
Let the velocity of the first body be V1,given that it is projected at an angle of 60
Let the velocity of the second body be V2,given that it is projected at an angle of 30 
The given condition here is that the maximum height attained by both bodies is the same i.e Hmax is same for both of them 
Let the maximum height attained by first body is Hmax-1 and maximum height attained by the second body is Hmax-2 
We know that Hmax is equal 
this implies that Hmax-1=Hmax-2 
V12sin2\alpha/2g=V22sin2\beta/2g
V12sin260=V22sin230
V12/ V22 =sin230/sin260
V1/V2=1/\sqrt3